OpenStudy (anonymous):

Solve 4x + 3y = 8x + y for y. Answer y = 2x y = 4x + 2 y = 8x - 4

OpenStudy (anonymous):

y = 2x

OpenStudy (anonymous):

y = 2x ----> 4x + 6x = 8x +2x

OpenStudy (anonymous):

Combine like terms 8x-4x=3y-y 4x=2y (divide by 2) y=2x

OpenStudy (anonymous):

Thanks guys!